# Benchmarking Cross-Domain Audio-Visual Deception Detection

### v2

@@ 4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
Following the implementation by Ganin et al. [49], we compared the multi-to-single domain generalization accuracies w/w.o GRL. GRL was proposed to mitigate the domain shift issue by manipulating the training gradients. It worked by acting as an identity transform in forward propagation and multiplying the gradient by a certain negative constant during the backpropagation without having trainable parameters. GRL was inserted between encoders and domain classifiers, which was easy to implement. As GRL is a widely adopted method for domain generalization, it is investigated to show its effectiveness for the deception detection task. We selected domain-simultaneous as the baseline and added GRL to the original network with the same training setups. The average accuracies were reported in Fig. 3, where different types of visual and audio features and methods were compared. Training with GRL, the performance of ResNet18 and ResNet18+GRU models using visual (face frames) features and Wave2Vec model using waveform were enhanced. However, we observed that MLP models using visual (AU/gaze/affect) features and the ResNet18 model using Mel spectrograms degraded in performance. Generally, ResNet18 trained with GRL performed better than MLP for visual modality, and Wave2Vec trained with GRL boosted the performance and surpassed the model trained on the Mel spectrogram.
@@
TABLE V: The fusion results of multi-to-single cross-domain generalization accuracy (%) for different generalization strategies.
Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
@@
Score-level Average 53.23 43.56 57.43 50.94 51.40 58.42 48.51 51.93
Concat 52.00 56.44 60.40 51.88 51.40 62.38 57.43 55.99
SE-Concat 55.69 58.42 44.55 52.50 56.07 58.42 58.42 54.87
Cross-Atten 51.08 52.48 55.45 51.56 57.01 60.40 57.43 55.06
MLP-Mixer 53.85 43.56 62.38 52.19 56.07 61.39 59.41 55.55
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 58.43
Score-level Average 49.54 55.45 52.48 51.25 54.21 59.41 55.45 53.97
Concat 49.54 54.46 49.50 53.75 44.86 63.37 54.46 52.85
SE-Concat 52.00 58.42 42.57 51.56 54.21 58.42 63.37 54.36
Cross-Atten 50.46 52.48 60.40 53.12 52.34 58.42 59.41 55.23
MLP-Mixer 53.54 57.43 52.48 50.31 52.34 60.40 47.52 53.43
@@
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 58.14
Score-level Average 48.62 58.42 56.44 50.94 51.40 56.44 53.47 53.68
Concat 48.31 53.47 56.44 50.31 56.07 63.37 57.43 55.06
SE-Concat 49.85 50.50 47.52 52.19 49.53 57.43 61.39 52.63
Cross-Atten 51.38 63.37 61.39 51.25 48.60 59.41 58.42 56.26
MLP-Mixer 55.69 59.41 57.43 53.75 50.47 61.39 50.50 55.52
@@
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 56.78
Score-level Average 54.77 57.43 57.43 52.81 60.75 58.42 56.44 56.86
Concat 54.46 54.46 47.52 53.12 53.27 61.39 56.44 54.38
SE-Concat 53.23 55.45 54.46 52.50 59.81 56.44 58.42 55.76
Cross-Atten 48.62 57.43 50.50 53.12 48.60 58.42 56.44 53.30
MLP-Mixer 48.92 48.51 59.41 52.81 54.21 57.43 51.49 53.25
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.62
@@
Score-level Average 55.38 57.43 58.42 50.94 57.94 50.50 61.39 56.00
Concat 55.08 58.42 63.37 51.56 62.62 62.38 53.47 58.13
SE-Concat 50.46 49.50 63.37 52.19 57.01 65.35 56.44 56.33
Cross-Atten 55.08 59.41 60.40 52.50 48.60 58.42 61.39 56.54
MLP-Mixer 57.54 63.37 63.37 50.94 65.42 52.48 56.44 58.51
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.62 61.39 63.37 51.56 58.88 62.38 58.42 58.95
Score-level Average 50.15 58.42 60.40 50.62 51.40 60.40 58.42 55.69
Concat 53.23 62.38 54.46 51.25 59.81 54.46 54.46 55.72
SE-Concat 50.15 52.48 64.36 50.31 59.81 55.45 50.50 54.72
Cross-Atten 52.92 57.43 49.50 52.81 65.42 53.47 66.34 56.84
MLP-Mixer 51.69 58.42 57.43 51.25 62.62 58.42 56.44 56.61
@@
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.23 63.37 56.44 51.25 63.55 59.41 56.44 57.67
Score-level Average 51.38 56.44 54.46 50.00 64.49 61.39 49.50 55.38
Concat 49.85 62.38 58.42 51.88 61.68 55.45 51.49 55.88
SE-Concat 50.46 62.38 58.42 50.00 57.94 59.41 65.35 57.71
Cross-Atten 49.85 62.38 50.50 53.44 58.88 56.44 56.44 55.42
MLP-Mixer 50.15 55.45 58.42 50.31 58.88 61.39 57.43 56.00
@@
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 50.15 59.41 50.50 52.19 58.88 60.40 61.39 56.13
Score-level Average 56.00 52.48 59.41 51.88 51.40 60.40 56.44 55.43
Concat 51.69 58.42 58.42 52.81 57.01 61.39 56.44 56.60
SE-Concat 56.31 59.41 58.42 50.31 53.27 61.39 55.45 56.37
Cross-Atten 49.54 67.33 54.46 51.56 66.36 52.48 60.40 57.45
MLP-Mixer 50.15 63.37 60.40 51.56 62.62 61.39 61.39 58.70
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 66.34 60.40 51.88 59.81 61.39 62.38 59.13
@@
Score-level Average 54.77 46.53 59.41 50.62 57.94 59.41 58.42 55.30
Concat 53.85 53.47 63.37 51.88 57.94 65.35 58.42 57.75
SE-Concat 56.62 44.55 63.37 52.19 54.21 58.42 59.41 55.54
Cross-Atten 54.15 49.50 62.38 54.69 57.01 60.40 63.37 57.36
MLP-Mixer 57.54 62.38 71.29 51.25 54.21 62.38 62.38 60.20

### v3

@@ 4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
Gradient Reversal Layer (GRL) [52] has become a foundational component in domain adaptation and domain generalization due to its ability to encourage domain-invariant feature learning. Its effectiveness and versatility have been demonstrated across diverse fields such as face anti-spoofing [42], sleep stage classification [53], multi-view 3D object detection [54], etc.
@@
Following the implementation by Ganin and Lempitsky [52], we compared the multi-to-single domain generalization accuracies w/w.o GRL. GRL was proposed to mitigate the domain shift issue by manipulating the training gradients. It worked by acting as an identity transform in forward propagation and multiplying the gradient by a certain negative constant during the backpropagation without having trainable parameters. GRL was inserted between encoders and domain classifiers, which was easy to implement. As GRL is a widely adopted method for domain generalization, it is investigated to show its effectiveness for the deception detection task. We selected domain-simultaneous as the baseline and added GRL to the original network with the same training setups. The average accuracies were reported in Fig. 3, where different types of visual and audio features and methods were compared. Training with GRL, the performance of ResNet18 and ResNet18+GRU models using visual (face frames) features and Wave2Vec model using waveform were enhanced. However, we observed that MLP models using visual (AU/gaze/affect) features and the ResNet18 model using Mel spectrograms degraded in performance. Generally, ResNet18 trained with GRL performed better than MLP for visual modality, and Wave2Vec trained with GRL boosted the performance and surpassed the model trained on the Mel spectrogram.
@@
TABLE V: Fusion results of single-to-single cross-domain generalization accuracy (%).
Modality & Inputs Fusion Postion Fusion Method R to B1 R to B2 R to M R to D R to E B1 to R B1 to B2 B1 to M B1 to D B1 to E
Score-level Average 53.23 41.58 51.88 53.85 50.16 64.49 62.38 50.62 52.29 50.97
Concat 51.08 55.45 51.88 52.90 50.49 54.21 58.42 51.25 51.94 50.08
SE-Concat 53.85 60.40 51.25 53.07 51.29 55.14 58.42 50.62 55.32 50.81
Cross-Atten 55.38 61.39 52.19 54.11 52.35 51.40 60.40 51.25 56.09 51.05
MLP-Mixer 55.08 48.51 53.44 54.02 52.18 56.07 58.42 53.75 57.22 51.29
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.92 59.41 57.94 55.32 53.22 63.37 53.75 53.75 57.39 51.38
Score-level Average 53.23 49.50 51.88 53.69 50.81 50.47 59.41 53.75 52.90 50.49
Concat 50.77 53.47 50.47 52.64 50.24 62.38 51.56 51.88 51.69 49.35
SE-Concat 50.15 44.55 51.40 51.86 50.16 61.39 53.12 52.50 54.02 50.24
Cross-Atten 54.46 51.49 55.14 53.93 51.21 58.42 51.25 52.19 54.80 50.49
MLP-Mixer 52.31 55.45 57.94 54.36 51.54 63.37 53.12 51.25 55.66 51.21
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.54 55.45 56.07 55.06 51.78 61.39 50.94 53.12 57.13 52.35
Score-level Average 49.85 58.42 54.06 52.16 50.08 45.79 60.40 50.62 51.08 50.57
Concat 49.54 53.47 47.66 50.91 50.57 61.39 53.44 51.88 50.39 50.49
SE-Concat 50.15 48.51 55.14 50.99 50.65 60.40 53.12 50.00 52.93 50.24
Cross-Atten 53.23 44.55 57.94 51.25 50.89 55.45 54.69 54.06 55.14 51.38
MLP-Mixer 49.54 57.43 50.47 53.33 51.38 63.37 54.06 53.12 56.18 51.70
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.52 54.46 57.94 54.97 51.13 61.39 53.12 51.56 57.48 52.43
Score-level Average 52.00 58.42 51.88 52.43 50.97 53.27 58.42 50.94 54.11 50.16
Concat 53.23 58.42 55.14 53.93 50.32 61.39 52.50 52.19 55.23 50.32
SE-Concat 51.38 58.42 51.40 53.59 50.24 62.38 52.81 50.94 56.53 51.62
Cross-Atten 51.08 48.51 55.14 54.11 51.38 60.40 53.12 53.44 57.22 50.08
MLP-Mixer 55.69 46.53 44.86 54.97 51.05 63.37 51.56 50.94 57.04 51.21
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 60.40 57.01 56.35 51.86 64.36 53.44 51.25 57.56 51.62
Modality & Inputs Fusion Postion Fusion Method M to R M to B1 M to B2 M to D M to E D to R D to B1 D to B2 D to M D to E
Score-level Average 57.94 53.23 62.38 53.41 50.32 52.31 51.69 52.28 50.46 50.32
Concat 57.01 50.15 61.39 53.50 51.21 53.84 50.46 54.08 50.40 50.16
SE-Concat 56.07 51.38 65.35 54.19 50.57 54.42 51.38 54.48 50.37 50.77
Cross-Atten 55.14 56.31 60.40 54.39 51.13 55.23 52.62 57.43 50.86 50.97
MLP-Mixer 55.14 59.08 59.41 55.14 51.29 55.08 52.92 59.41 50.77 51.05
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 60.75 56.00 61.39 54.97 52.02 56.92 52.10 60.04 51.49 51.69
Score-level Average 65.42 56.31 49.50 52.90 50.49 52.00 51.08 53.47 50.24 50.18
Concat 54.21 52.62 58.42 51.60 50.81 54.15 50.77 55.45 50.47 50.52
SE-Concat 65.42 56.31 66.34 52.03 51.05 54.46 50.96 56.64 50.60 50.81
Cross-Atten 63.55 56.62 66.34 52.38 51.29 54.46 50.50 56.84 51.20 51.08
MLP-Mixer 64.49 57.85 62.38 52.20 51.05 56.00 50.46 57.43 51.08 51.38
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 67.29 57.85 64.36 55.83 52.35 56.31 51.69 58.42 53.27 52.30
Score-level Average 49.53 56.00 63.37 52.03 51.38 53.38 50.35 53.26 51.40 50.89
Concat 57.01 50.15 58.42 52.29 50.97 56.25 50.15 53.47 52.34 51.05
SE-Concat 57.01 49.85 63.37 52.38 50.57 56.62 50.77 54.46 53.64 51.29
Cross-Atten 63.55 51.69 64.36 53.07 51.05 56.31 51.69 55.35 52.92 51.78
MLP-Mixer 59.81 55.08 69.31 54.36 56.85 53.12 51.38 55.42 52.62 52.67
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 69.16 58.15 64.36 55.40 51.46 57.23 53.85 61.39 54.21 53.32
Score-level Average 57.01 53.54 61.39 50.82 51.38 52.92 52.31 58.22 51.60 52.04
Concat 56.07 54.15 60.40 51.69 50.16 53.23 52.00 58.32 51.40 52.00
SE-Concat 59.81 54.77 52.48 52.03 51.78 56.95 53.65 59.45 51.50 52.40
Cross-Atten 60.75 56.31 60.40 53.24 51.94 57.33 54.15 59.41 52.30 52.34
MLP-Mixer 64.49 56.00 60.40 54.11 51.54 57.54 54.46 60.42 52.31 53.27
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 67.29 56.00 62.38 55.32 52.91 58.77 55.08 63.37 53.85 53.54
Modality & Inputs Fusion Postion Fusion Method E to R E to B1 E to B2 E to M E to D Avg
Score-level Average 49.68 51.21 56.46 49.53 53.50 53.05
Concat 50.73 51.49 56.33 50.10 52.38 52.84
SE-Concat 50.16 52.48 57.43 50.25 53.24 53.71
Cross-Atten 50.65 53.46 57.83 50.47 53.07 54.22
MLP-Mixer 51.38 55.42 57.25 51.24 54.11 54.35
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 52.27 56.21 60.09 51.45 54.54 55.78
Score-level Average 49.84 52.46 57.87 50.03 52.72 52.83
Concat 51.21 52.69 58.72 50.33 51.34 52.71
SE-Concat 50.97 53.62 59.41 50.54 52.81 53.65
Cross-Atten 50.57 54.99 59.70 50.67 52.72 54.25
MLP-Mixer 51.21 55.08 60.08 50.95 52.20 54.80
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.70 55.45 61.39 51.63 54.19 55.79
Score-level Average 50.24 54.50 57.43 50.65 51.08 52.74
Concat 50.32 54.46 58.42 51.70 51.86 52.74
SE-Concat 50.81 54.66 58.22 51.38 51.96 53.16
Cross-Atten 50.97 55.32 58.62 52.18 53.85 54.05
MLP-Mixer 50.65 55.65 58.77 52.10 54.11 54.90
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.02 56.44 60.41 53.16 55.06 56.14
Score-level Average 49.50 53.23 58.00 51.21 52.29 53.52
Concat 50.50 53.47 58.92 51.13 52.90 53.96
SE-Concat 50.47 54.27 59.11 51.29 54.02 54.13
Cross-Atten 50.67 55.45 61.29 52.02 54.28 54.65
MLP-Mixer 51.49 56.31 60.40 52.27 55.92 54.73
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.48 56.25 63.37 54.91 56.09 56.82
TABLE VI: Fusion results for multi-to-single cross-domain accuracy (%) using domain-simultaneous sampling.
@@
Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&M to D R&M to E R&B1 to B2 R&B1 to M R&B1 to D R&B1 to E B1&M to R B1&M to B2 R&B1&M to B2
Score-level Average 53.23 43.56 50.57 50.16 57.43 50.94 56.44 50.16 51.40 58.42 48.51
Concat 52.00 56.44 51.29 50.86 60.40 51.88 57.04 51.70 51.40 62.38 57.43
SE-Concat 55.69 58.42 52.08 52.84 44.55 52.50 56.87 51.13 56.07 58.42 58.42
Cross-Atten 51.08 52.48 52.51 52.24 55.45 51.56 57.13 51.94 57.01 60.40 57.43
MLP-Mixer 53.85 43.56 53.02 53.32 62.38 52.19 57.74 52.10 56.07 61.39 59.41
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.23 57.43 53.50 53.40 63.37 52.19 57.39 51.86 57.01 63.37 62.38
Score-level Average 49.54 55.45 50.73 50.73 52.48 51.25 57.56 50.81 54.21 59.41 55.45
Concat 49.54 54.46 50.97 50.92 49.50 53.75 56.96 51.05 44.86 63.37 54.46
SE-Concat 52.00 58.42 51.54 52.00 42.57 51.56 57.22 51.46 54.21 58.42 63.37
Cross-Atten 50.46 52.48 52.02 53.08 60.40 53.12 57.30 51.13 52.34 58.42 59.41
MLP-Mixer 53.54 57.43 52.10 53.32 52.48 50.31 57.56 52.18 52.34 60.40 47.52
@@
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 64.36 53.97 53.43 56.44 53.75 57.82 52.35 58.88 59.41 58.42
Score-level Average 48.62 58.42 51.21 50.29 56.44 50.94 56.48 51.78 51.40 56.44 53.47
Concat 48.31 53.47 51.86 50.44 56.44 50.31 57.13 51.29 56.07 63.37 57.43
SE-Concat 49.85 50.50 53.87 52.32 47.52 52.19 57.42 52.02 49.53 57.43 61.39
Cross-Atten 51.38 63.37 54.50 53.24 61.39 51.25 57.74 52.27 48.60 59.41 58.42
MLP-Mixer 55.69 59.41 55.06 53.24 57.43 53.75 57.43 52.99 50.47 61.39 50.50
@@
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 55.45 57.39 54.20 58.42 54.37 58.02 52.83 52.34 61.39 60.40
Score-level Average 54.77 57.43 55.50 50.90 57.43 52.81 55.45 51.46 60.75 58.42 56.44
Concat 54.46 54.46 55.52 50.82 47.52 53.12 55.85 51.78 53.27 61.39 56.44
SE-Concat 53.23 55.45 56.00 51.60 54.46 52.50 56.33 51.05 59.81 56.44 58.42
Cross-Atten 48.62 57.43 56.26 52.18 50.50 53.12 57.46 51.62 48.60 58.42 56.44
MLP-Mixer 48.92 48.51 56.61 53.68 59.41 52.81 58.40 52.21 54.21 57.43 51.49
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.31 59.41 57.30 54.57 58.42 52.50 59.41 52.84 58.88 61.39 60.40
Modality & Inputs Fusion Postion Fusion Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
Score-level Average 50.66 53.56 52.50 52.34 49.96 50.47 52.29 50.44 53.24 51.19 51.18
Concat 51.38 54.46 50.31 52.31 53.33 52.18 53.40 51.25 53.24 51.19 51.18
SE-Concat 51.49 54.86 51.85 52.20 55.49 53.07 55.15 51.88 54.87 54.81 53.94
Cross-Atten 52.00 55.45 52.06 53.20 56.70 54.96 55.45 52.31 55.31 56.56 54.44
MLP-Mixer 52.48 56.42 52.94 53.27 56.96 56.44 56.44 53.13 57.24 57.98 55.16
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.47 59.01 53.62 54.21 58.43 57.10 57.54 53.75 60.40 60.99 56.84
Score-level Average 50.50 53.47 51.25 52.64 54.80 50.21 52.47 50.62 53.94 55.00 52.98
Concat 52.34 53.87 50.62 52.98 56.78 53.48 54.26 51.55 53.69 56.16 53.21
SE-Concat 52.10 55.15 51.38 53.85 58.08 55.75 54.35 53.25 54.06 56.26 54.14
Cross-Atten 53.84 56.44 52.94 54.61 58.17 57.39 55.20 53.69 55.13 57.15 54.99
MLP-Mixer 53.18 56.74 53.38 54.35 58.30 58.42 57.37 54.06 59.94 58.22 54.91
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.24 59.41 53.50 55.14 58.53 58.65 58.92 54.88 62.38 59.35 57.12
Score-level Average 52.16 56.44 51.81 53.50 52.72 50.14 57.00 50.94 53.62 54.98 53.28
Concat 52.08 56.04 51.20 53.44 56.87 50.77 57.50 51.57 54.69 57.39 54.18
SE-Concat 53.21 57.27 52.75 54.36 57.13 56.09 58.51 52.69 55.75 58.22 54.29
Cross-Atten 53.85 59.64 52.38 54.45 57.74 56.44 58.49 53.38 57.25 59.61 55.93
MLP-Mixer 54.14 60.40 53.56 55.64 58.25 57.03 59.41 54.50 59.38 59.04 56.13
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.11 62.38 54.02 56.28 58.48 57.65 60.19 54.64 60.50 60.13 57.11
Score-level Average 53.85 54.46 52.69 53.21 56.27 50.84 54.13 50.93 53.13 57.04 54.66
Concat 55.42 57.43 52.62 54.97 56.61 53.33 54.87 51.50 54.06 57.54 54.43
SE-Concat 56.44 58.13 52.90 56.07 56.88 56.18 55.02 53.16 54.44 57.22 55.32
Cross-Atten 57.43 59.41 53.20 56.21 58.22 57.22 57.98 53.29 56.93 58.13 55.17
MLP-Mixer 56.46 60.34 54.00 57.07 58.42 58.08 59.63 54.31 59.75 59.44 55.77
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.48 63.17 54.35 58.88 59.15 58.96 60.42 54.56 62.58 60.42 57.97
TABLE VII: Fusion results for multi-to-single cross-domain accuracy (%) using domain-alternating sampling.
@@
Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&M to D R&M to E R&B1 to B2 R&B1 to M R&B1 to D R&B1 to E B1&M to R B1&M to B2 R&B1&M to B2
Score-level Average 55.38 57.43 50.67 50.32 58.42 50.94 55.32 50.22 57.94 50.50 61.39
Concat 55.08 58.42 50.29 50.54 63.37 51.56 56.98 52.73 62.62 62.38 53.47
SE-Concat 50.46 49.50 51.18 52.78 63.37 52.19 56.80 52.05 57.01 65.35 56.44
Cross-Atten 55.08 59.41 52.41 53.01 60.40 52.50 57.71 52.13 48.60 58.42 61.39
MLP-Mixer 57.54 63.37 53.23 53.22 63.37 50.94 57.82 52.24 65.42 52.48 56.44
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.62 61.39 53.60 53.50 63.37 51.56 57.30 52.65 58.88 62.38 58.42
Score-level Average 50.15 58.42 50.65 50.87 60.40 50.62 57.88 50.06 51.40 60.40 58.42
Concat 53.23 62.38 51.11 50.95 54.46 51.25 57.04 51.24 59.81 54.46 54.46
SE-Concat 50.15 52.48 51.66 52.55 64.36 50.31 57.73 51.26 59.81 55.45 50.50
Cross-Atten 52.92 57.43 52.13 53.89 49.50 52.81 57.82 52.08 65.42 53.47 66.34
MLP-Mixer 51.69 58.42 52.15 53.99 57.43 51.25 57.68 52.11 62.62 58.42 56.44
@@
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.23 63.37 54.01 53.53 56.44 51.25 57.85 52.32 63.55 59.41 56.44
Score-level Average 51.38 56.44 51.20 50.38 54.46 50.00 56.05 51.74 64.49 61.39 49.50
Concat 49.85 62.38 51.00 50.48 58.42 51.88 57.23 51.86 61.68 55.45 51.49
SE-Concat 50.46 62.38 53.24 53.84 58.42 50.00 57.72 52.34 57.94 59.41 65.35
Cross-Atten 49.85 62.38 54.51 53.92 50.50 53.44 57.81 52.88 58.88 56.44 56.44
MLP-Mixer 50.15 55.45 54.59 53.49 58.42 50.31 58.03 53.01 58.88 61.39 57.43
@@
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 50.15 59.41 54.99 54.31 50.50 52.19 58.25 53.06 58.88 60.40 61.39
Score-level Average 56.00 52.48 55.22 50.92 59.41 51.88 55.06 50.16 51.40 60.40 56.44
Concat 51.69 58.42 55.38 51.09 58.42 52.81 55.32 51.25 57.01 61.39 56.44
SE-Concat 56.31 59.41 56.22 51.28 58.42 50.31 56.38 51.99 53.27 61.39 55.45
Cross-Atten 49.54 67.33 56.43 52.39 54.46 51.56 57.06 52.89 66.36 52.48 60.40
MLP-Mixer 50.15 63.37 57.00 53.33 60.40 51.56 58.03 53.09 62.62 61.39 61.39
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 66.34 57.38 55.02 60.40 51.88 59.82 53.86 59.81 61.39 62.38
Modality & Inputs Fusion Postion Fusion Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
Score-level Average 50.65 53.45 52.35 51.16 48.96 50.67 52.31 50.35 54.34 52.34 53.10
Concat 51.42 54.44 52.33 51.34 54.32 52.31 53.31 51.29 54.11 54.39 54.60
SE-Concat 51.48 54.38 52.06 52.45 56.03 53.28 54.29 51.85 54.65 54.82 54.40
Cross-Atten 52.34 55.26 53.10 53.52 56.13 54.65 54.99 53.05 55.33 55.69 55.01
MLP-Mixer 52.46 55.39 53.15 53.65 56.35 56.33 56.05 53.38 56.89 57.82 56.07
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.89 57.02 53.45 54.62 58.16 57.01 58.24 54.02 59.89 59.85 56.94
Score-level Average 50.53 52.88 51.38 52.63 54.28 51.04 52.85 51.25 53.22 55.68 53.57
Concat 51.45 53.66 52.36 52.89 56.35 53.28 54.39 52.22 53.31 56.83 54.15
SE-Concat 52.11 55.02 52.39 53.08 58.04 54.99 54.78 53.48 54.09 56.89 54.34
Cross-Atten 52.98 56.36 53.24 53.96 58.26 56.84 55.76 53.87 55.68 57.63 55.62
MLP-Mixer 53.42 56.79 53.33 54.21 58.45 57.05 57.35 54.33 59.88 58.47 55.98
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.64 58.98 53.06 54.76 58.95 58.28 58.92 55.21 62.05 59.86 56.96
Score-level Average 52.17 56.42 51.28 52.36 53.15 50.58 57.08 51.43 53.34 54.27 53.79
Concat 52.89 56.63 51.64 53.34 56.24 50.87 57.06 52.11 54.99 57.25 54.50
SE-Concat 53.42 57.06 52.03 54.68 57.03 56.92 58.45 52.98 55.56 58.62 56.07
Cross-Atten 53.22 58.99 52.65 54.92 57.05 56.83 58.69 53.89 57.84 59.37 55.74
MLP-Mixer 54.65 59.78 53.08 55.01 58.34 57.36 59.05 54.54 59.08 59.38 56.26
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.38 61.49 54.56 56.38 58.65 57.82 60.18 55.38 61.05 61.20 56.89
Score-level Average 52.85 56.84 52.41 53.04 56.48 51.22 54.28 51.02 53.45 57.39 54.21
Concat 54.86 56.98 52.38 54.02 56.87 53.23 54.62 52.24 54.89 57.26 55.07
SE-Concat 55.53 57.32 53.06 55.93 56.92 56.35 55.23 53.26 56.53 58.02 55.65
Cross-Atten 56.88 59.63 53.66 55.89 57.98 57.26 57.01 53.82 56.87 58.25 56.58
MLP-Mixer 56.08 61.02 55.24 57.46 58.35 58.35 59.55 54.38 57.05 59.36 57.58
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.23 63.93 55.05 58.95 59.65 59.01 61.14 55.01 59.96 61.35 58.63
TABLE VIII: Fusion results for multi-to-single cross-domain accuracy (%) using domain-by-domain sampling.
@@
Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&M to D R&M to E R&B1 to B2 R&B1 to M R&B1 to D R&B1 to E B1&M to R B1&M to B2 R&B1&M to B2
Score-level Average 54.77 46.53 50.73 50.45 59.41 50.62 55.68 50.43 57.94 59.41 58.42
Concat 53.85 53.47 50.37 51.02 63.37 51.88 56.32 52.39 57.94 65.35 58.42
SE-Concat 56.62 44.55 51.23 52.39 63.37 52.19 56.56 53.56 54.21 58.42 59.41
Cross-Atten 54.15 49.50 52.46 53.34 62.38 54.69 56.98 53.21 57.01 60.40 63.37
MLP-Mixer 57.54 62.38 53.53 53.45 71.29 51.25 57.59 53.49 54.21 62.38 62.38